OpenStudy (anonymous):

Given f(x) = 3x + 2 and g(x) = x2 + 4, find (g • f )(2).

OpenStudy (anonymous):

do know what to do at all?

OpenStudy (anonymous):

ok, so what you do is plug in the 2 for the value x.

OpenStudy (anonymous):

you do this for both equations; f(x) and g(x).

OpenStudy (anonymous):

then you multiply the results from both.

OpenStudy (anonymous):

Oh, and btw, welcome to openstudy! :D

OpenStudy (anonymous):

so for f(x): f(x) = 3x + 2 f(x) = 3(2) + 2 f(x) = 6 + 2 f(x) = 8 for g(x): g(x) = x^2 + 4 g(x) = (2)^2 + 4 g(x) = 4 + 4 g(x) = 8 then you multiply them together: 8*8 = your solution

OpenStudy (anonymous):

hope this helps :D

hero (hero):

(g • f )(x) = (3x + 2)(x^2 + 4) = x^2(3x + 2) + 4(3x + 2) = 3x^3 + 2x^2 + 12x + 8 (g • f )(2) = 3(2)^3 + 2(2)^2 + 12(2) + 8 = 3(8) + 2(4) + 24 + 8 = 24 + 8 + 24 + 8 = 24 + 24 + 8 + 8 = 48 + 16 = 64

OpenStudy (anonymous):

that method works too... which is more conventional @Hero?

hero (hero):

Well basically as a rule, (g • f )(x) = g(x)f(x) The question said to calculate (g • f )(2) You computed g(2)*f(2) They are both equivalent computations. One method seems "easier" to compute than the other.

OpenStudy (anonymous):

ok :D thanks!

OpenStudy (anonymous):

oh... i feel like a idiot. Ive done this before in class. I didn't recognize it here though. what is this called btw.

hero (hero):

Operations on Functions: (f + g)(x) = f(x) + g(x) (f - g)(x) = f(x) - g(x) (f • g)(x) = f(x) • g(x) (f ÷ g)(x) = f(x) ÷ g(x)